A sample of O2 gas with a volume of 8.22 L was collected at 199 °C. What is the volume of the gas if the temperature is changed to -31 °C at constant pressure?
1
Expert's answer
2017-03-06T13:35:05-0500
Charle’s law:
V_1/T_1 =V_2/T_2 =const
V_2=(V_1×T_2)/T_1 =(8.22L×242.15K)/472.15K=4.22L

Answer: V=4.22L